# HumanVBench: Exploring Human-Centric Video Understanding Capabilities of MLLMs with Synthetic Benchmark Data

### v2

@@ 5 Conclusion
Our work opens several avenues for future research. These include expanding the benchmark to encompass broader scenarios, such as video generation tasks; enhancing model architectures to improve temporal and multimodal fusion; and refining datasets to more effectively capture the nuances of human emotions and intentions. By open-sourcing the benchmark and underlying methodologies, we hope HumanVBench can foster collaborative efforts aimed at advancing the frontiers of human-like video understanding capabilities in MLLMs.
@@
6 Overview
In the appendix, we first We provide more benchmark statistics in Section 7, then the modality ablation experiments in VideoLLaMA2 in Section 8, followed by additional evaluation details in Section 9, followed by the detailed definition and examples for the 16 tasks of HumanVBench in Section 10. Then, we present implementation specifics of each operator for the proposed Human-Centric Annotation Pipeline, illustrated by an example of the annotation process in Section 11. Finally, we introduce the construction details of all tasks and the work of the human annotators in Section 12.
7 More Statistics of HumanVBench
HumanVBench focuses on short video understanding, specifically videos with a duration of 10 seconds or less. It includes a total of 2116 question instances, with the specific number for each task indicated in 1. The total video duration amounts to 4.7 hours and demonstrates a variety of people, scenes, and video shooting styles, as shown in 6.
@@ 7 More Statistics of HumanVBench
8 Modality Ablation in VideoLLaMA2
@@
Models Input Modal Human Emotion Perception Person Recognition Human Behavior Analysis Speech-Visual Alignment
ER ETA AR EIC Avg T2H H2T HC ATD Avg BTA BCA AST TSA Avg AVSM ASD AVAD SCM Avg
Random 25 25.0 25.0 25.0 22.9 27.9 25.0 23.1 25.0 25.3 25.0 25.0 25.0 20.0 23.8 42.8 23.6 33.3 25.0 31.2
Video-LLaMA2.1-AV A, V 32.3 28.7 47.0 21.7 32.4 34.6 47.9 41.5 16.7 35.2 47.5 43.3 27.0 33.3 37.8 44.0 31.6 32.1 23.7 32.9
Video-LLaMA2.1-AV V 27.1 31.6 45.1 17.2 30.3 36.9 46.8 36.8 15.0 33.9 43.2 50.7 31.0 29.9 38.7 43.4 29 32.8 18.8 31.0
VideoLLaMA2-7B-16F V 32.3 30.9 34.8 23.7 30.4 37.4 51.1 36.8 15.8 35.3 53.4 50.7 32.0 49.2 46.3 47.4 38.1 32.1 15.4 33.3
@@
9 Model Evaluation Implementation
Prompt. In order to facilitate the statistical model to answer the results, following common practices used in MLLM evaluations [24, 18], we adopt the following prompt to guide the MLLM to output option letters: “ Select the best answer to the following multiple-choice question based on the video. Respond with only the letter of the correct option. <Question-choices> Only answer best answer’s option letter. Your option is: ”. Evaluation Environments. All evaluation experiments for open-source models were conducted on a single NVIDIA L20 GPU with an inference batch size of 1.
@@ 9 Model Evaluation Implementation
Model Time (min) top_p top_k num_beams temp. VRAM
Chat-UniVi (7B) 35 1 50 1 0.2 14G
CogVLM2-Video (8B) 34 0.1 1 1 0.2 26G
Video-LLaVA (7B) 46 1 50 1 1 37G
LLaVA-OneVision (7B) 46 1 50 1 1 33G
PLLaVA (7B) 35 0.9 50 1 1 18G
ShareGPT4Video (8B) 44 0.9 50 1 1 17G
Otter-V (7B) 60 1 50 3 1 17G
VideoChat2-IT (7B) 50 0.9 50 1 1 18G
InternVL2 (7B) 44 1 50 1 1.0 21G
InternVL2.5 (7B) 31 1 50 1 1.0 21G
Qwen2-VL (7B) 43 0.001 1 1 0.1 22G
Qwen2.5-VL (7B) 33 0.001 1 1 0.1 22G
LLaVA-Video (7B) 173 0.8 20 1 0.7 36G
Video-LLaMA3 (7B) 86 0.8 20 1 0.7 21G
Video-LLaMA (7B) 38 1 50 2 1 22G
Video-LLaMA2.1 (7B) 27 0.9 50 1 0.2 23G
ImageBind-LLM (7B) 75 1 50 1 1 22G
ChatBridge (13B) 27 1 50 1 0.2 28G
OneLLM (7B) 116 0.75 50 1 0.1 16G
GPT-4o 185 1 - - 1 API
Gemini-1.5-Pro 250 1 40 - 0.9 API
@@
10 Definitions and Examples for Each Task
10.1 Emotion Perception
@@
Figure 10: Example of Emotion Intensity Compare task.
10.2 Person Recognition
@@
10.3 Human Behavior Analysis
@@
10.4 Cross-Modal Speech-Visual Alignment
@@
11 Annotations Details and Examples in Human-Centric Annotation Pipeline
@@
In the video_facial_description_mapper, we use the face bounding box track to crop the video, creating face-focused reconstructed videos for emotion description using VideoLLaMA2.1 [16]. The choice of VideoLLaMA2.1 is based on a comparative analysis of multiple models, which revealed that VideoLLaMA2.1 is often more effective at identifying negative emotions. This capability is particularly important for adjusting emotion distribution before designing emotion recognition tasks.
@@
12 Complete Construction Details of All Tasks
@@
12.1 Construction Details of 6 Descriptive Human-Centric Questions
For these six tasks, the video-MLLM used to obtain task-oriented captions is VideoLLaMA-2.1[16]. The LLM used for generating question and initial answer pairs is Qwen2.5[54]. The three video-MLLMs employed for optimizing answers and producing raw distractors are VideoLLaMA-2.1[16], CogVLM2[21], and LLaVA-OneVision[28], respectively. The LLM responsible for generating distractors is Qwen2.5. We first present the general instruction templates in the six task generation processes.
The prompt template for the three Video-MLLMs used to refine answers and generate raw distractors is:
Please focus on the people whose heads are highlighted with red bounding boxes in the video and answer my question: ⟨Question⟩; Provide a brief response in one sentence.
The instruction used to compare the answers is:
Based on the video and my question: ⟨Question⟩Tell me which answer is better: (A). ⟨current model’s answer. ⟩(B). ⟨previous best answer⟩. Just answer (A) or (B).
@@ 12.1 Construction Details of 6 Descriptive Human-Centric Questions
12.2 Construction Details of 10 Closed-Ended Human-Centric Questions
@@
12.3 Details of Human Efforts for HumanVBench
@@
Manual Verification and Correction: Except for the tasks that are already reliable enough, which include: Emotion Recognition in Conversation task reconstructed from the MELD benchmark, the three tasks derived from the aforementioned human annotations, and the Audio-Visual Alignment Detection and Speech Content Matching tasks, all other tasks require manual verification to ensure quality, following the process described in Section 3.3.4. During correction, low-quality samples (e.g., person transitions in human tracking, video freezing midway) are required to be flagged for removal.
@@
Human-Evaluation Baseline in Table 1: For each task except the Emotion Recognition in Conversation task, we define the number of correct human responses based on the cross-checking process as follows: the sum of the number of questions answered correctly by the first annotator after completing all questions and the number of questions answered correctly by the second annotator on the disputed questions after correction by the first annotator. The human response accuracy for the task is calculated as the ratio of the number of correct human responses to the total number of questions in the task.

### v3

@@ 5 Conclusion
Beyond the benchmark itself, the core contribution of this work lies in its extensibility and efficient methodology. Our framework transforms model-generated errors into high-quality distractors and shifts the human role from creating tasks from scratch to efficiently validating and correcting them, thereby multiplying the benchmark’s production efficiency. By open-sourcing the benchmark and underlying methodologies, we hope HumanVBench can foster collaborative efforts aimed at advancing the frontiers of human-like video understanding capabilities in MLLMs.
Acknowledgments
This work was supported in part by the New Generation Artificial Intelligence-National Science and Technology Major Project (2025ZD0123003), the National Natural Science Foundation of China Enterprise Innovation and Development Joint Fund (Artificial Intelligence Field) Key Support Projects (U25B2072), and The Major Key Project of PCL (Grant No. PCL2025A17).
@@ Acknowledgments
Appendix Contents
  1. A.
    The Use of Large Language Models ....................................................................................................................A
  2. B.
    Comparison with Existing Video Benchmarks ....................................................................................................................B
  3. C.
    More Statistics of HumanVBench ....................................................................................................................C
  4. D.
    Full Results for 24 Models on 16 Tasks ....................................................................................................................D
  5. E.
    Distractor Quality Analysis ....................................................................................................................E
  6. F.
    Modality Ablation in VideoLLaMA2 ....................................................................................................................F
  7. G.
    Impact of Frame Sampling ....................................................................................................................G
  8. H.
    Impact of Human Density on Model Performance ....................................................................................................................H
  9. I.
    Model Evaluation Implementation ....................................................................................................................I
  10. J.
    Definitions and Examples for Each Task ....................................................................................................................J
  11. K.
    Annotations Details in Human-Centric Annotation Pipeline ....................................................................................................................K
  12. L.
    Complete Construction Details of All Tasks ....................................................................................................................L
Appendix A The Use of Large Language Models
The core research idea, including the conceptualization of HUMANVBENCH and its synthesis framework, originated entirely from the authors. Multimodal Large Language Models (MLLMs) and Large Language Models (LLMs) were then integrally utilized as a methodological component to efficiently automate data annotation and challenging question construction for the benchmark, a key focus of our research. Details of MLLM utilization, including the models, prompts, and annotation pipeline, are provided in the methodology section of the main paper (Section 3) and further elaborated in Appendix K and Appendix L. Additionally, LLMs assisted with language refinement during manuscript preparation. The authors assume full responsibility for all content, including research ideas, methodologies, and findings.
Appendix B Comparison with Existing Video Benchmarks
Table 3: Comparison of HumanVBench with existing video-based MLLM benchmarks.
Benchmark Feature Video #QA Human- centric QA Video Length Annotation Source QA Generation Evaluation Content
Emotion Person Recog. Behavior Lip-sync
Video-MME General Video
Comprehension
900 2,700 1,084
(40.1%)
<<1h Fully human
annotated
33 cases
(1%)
MVBench General Video
Comprehension
3,259 3,600 1,684
(46.7%)
<<35s Existing Captions LLM-based 33 cases
(0.9%)
121 cases
(3%)
ActivityNet-QA (Test Set) General Video
Comprehension
8,000 1,050
(13.1%)
$\sim$3min Fully human
annotated
TLQA Temporal Logic
Reasoning
160k <<5min Existing temporal labels Automatic logic templates
ANetQA Compositional
Reasoning
13.4M $\sim$3min Manually annotated scene graphs Automatic template-based
VEATIC Continuous Emotion (Valence/Arousal) 124 100% <<2m37s Real-time human labeling -
TVQA (Test Set) TV Plot
Understanding
1,809 7,623 100% <<90s Fully human annotated
Social-IQ Social Situations Reasoning 1,250 7,500 100% <<30s Fully human annotated
HumanVBench Foundational
Human Perception
2,182 2,475 100% <<30s Multi-operator auto-annotation Rule-based & LLM ensemble
Table 3 compares existing video benchmarks for MLLMs, highlighting their limited focus on foundational human perception tasks such as emotion understanding and audio-visual alignment. Moreover, most rely on manual annotations or repurposed datasets, which restricts scalability and applicability to unconstrained, in-the-wild videos In contrast, HumanVBench provides an automated, multi-operator pipeline and multi-model QA synthesis to deliver a comprehensive benchmark across key human-centric perceptual skills, enabling more precise and scalable evaluation of foundational human video understanding.
Appendix C More Statistics of HumanVBench
HumanVBench focuses on short video understanding, specifically videos with a duration of 30-second or less. It includes a total of 2475 question instances, with the specific number for each task indicated in 1. The total video duration amounts to 5.2 hours and demonstrates a variety of people, scenes, and video shooting styles, as shown in 6.
@@ Appendix C More Statistics of HumanVBench
Appendix D Full Results for 24 Models on 16 Tasks
We meticulously select 22 SOTA open-source video MLLMs. These included both visual-only models such as Qwen-VL series [3, 48], InternVL series [13, 12], VideoLLaMA3, InternVideo2.5 [49], ViLAMP [14],LLaVA-Video [63], ShareGPT4Video, CogVLM2-video [21], LLaVA-One-Vision [31], Chat-UniVi [25], VILA [37], VideoChat [33], and audio-visual models capable of analyzing both visual and audio inputs, such as the Qwen2.5-Omni [52], MIO [50], Video-LLaMA series [59, 15] and generalist MLLMs like ImageBind-LLM [18], ChatBridge [64], and OneLLM [19].
Table 4: Results on HumanVBench with 18 visual-only MLLMs, 7 audio-visual MLLMs, and 5 proprietary models. “Random” denotes random guessing, and “Human” indicates human-level performance. Task acronyms are defined in Figure 1. For each task, the best overall results are shown in bold, while the best open-source model results are underlined. “-” means the model recognizes its lack of required capabilities for the task and thus refuses to answer. The results for all open-source models are averaged over five runs with different random seeds.
@@ Appendix D Full Results for 24 Models on 16 Tasks
Emotion Perception Person Recognition Behavior Analysis Speech-Visual Alignment
Models ER ETA AR EIC Avg T2H H2T HC ATD Avg BTA BCA AST TSA Avg AVSM ASD AVAD SCM Avg
Random 24.6 24.7 25.0 23.1 24.4 27.9 24.6 23.1 25.0 25.2 23.0 23.6 25.0 20.0 22.9 42.8 23.6 33.3 25.0 31.2
ViLA 32.7 24.8 28.7 21.4 26.9 50.4 27.0 40.4 20.2 34.5 33.0 48.6 33.0 51.4 41.5 47.4 23.9 34.3 18.6 31.1
Video-LLaVA 18.3 18.8 26.2 8.0 17.8 27.9 40.6 28.3 31.7 32.1 30.8 32.5 27.0 40.7 32.8 50.0 28.4 34.3 21.2 33.5
Chat-Univi 26.4 15.0 10.9 16.5 17.2 29.1 27.0 17.8 19.8 23.4 27.9 39.4 14.0 6.7 22.0 42.8 20.6 26.3 18.6 27.1
VideoChat2-IT 33.5 29.1 39.6 26.8 32.3 20.1 47.3 11.2 26.7 26.3 45.8 43.7 34.0 24.9 37.1 43.4 27.7 31.4 23.0 31.4
InternVL2 36.1 31.3 40.2 30.4 34.5 70.4 61.2 37.2 20.8 47.4 49.8 52.3 38.0 33.8 43.5 51.7 55.0 33.6 25.5 41.5
Qwen-VL2 38.0 35.1 42.7 37.5 38.3 79.3 72.7 43.4 20.8 54.1 47.8 55.6 32.0 51.4 46.7 50.7 56.1 31.4 23.7 40.5
CogVLM2-Video 38.4 29.9 36.0 25.0 32.3 42.5 51.5 31.1 40.0 41.3 41.9 50.3 34.0 14.1 35.1 59.2 38.7 30.7 17.9 36.6
VideoLLaMA3 34.6 29.9 48.0 46.4 39.7 90.3 67.9 44.9 71.0 68.5 41.5 56.3 56.4 69.0 55.8 64.6 63.1 34.3 17.9 45.0
LLaVAOneVision 36.9 31.3 62.8 28.6 39.9 67.0 61.2 49.1 10.0 46.8 45.5 55.6 37.0 47.5 46.4 52.6 51.6 35.0 26.9 41.5
InternVideo 35.7 42.9 47.6 35.7 40.5 64.2 69.1 50.9 42.5 56.7 56.1 58.3 42.0 55.9 53.1 59.4 59.4 32.8 31.6 41.5
InternVL2.5 37.3 38.1 54.3 37.5 41.8 81.0 73.3 40.6 35.8 57.7 53.0 57.6 41.0 62.1 53.4 65.1 61.3 32.1 15.4 43.5
InternVL3 38.0 34.6 61.6 35.7 42.5 83.8 74.5 46.2 45.8 62.6 53.8 60.9 54.0 61.0 57.4 63.8 58.7 34.3 26.5 45.8
InternVL3.5 43.7 40.6 61.0 44.6 47.5 87.2 79.4 46.2 16.7 57.4 66.4 61.6 49.0 58.2 58.8 63.8 63.2 31.4 23.1 45.4
Qwen2.5-VL 43.0 30.6 35.4 50.0 39.8 88.8 74.5 50.9 30.8 61.3 51.0 58.3 34.0 50.8 48.5 71.1 61.3 33.6 18.8 46.2
Qwen3-VL 40.3 37.6 53.0 42.0 43.2 94.4 77.6 51.9 46.7 67.6 60.5 64.2 39.0 53.7 54.3 71.1 65.8 34.3 22.2 48.3
ShareGPT4Video 35.0 39.1 40.9 29.5 36.1 33.0 38.8 24.5 43.3 34.9 32.8 43.7 39.0 16.9 33.1 44.1 31.0 34.3 25.0 33.6
LLaVA-Video 35.4 32.3 59.8 28.6 39.0 74.3 61.8 44.3 26.7 51.8 41.9 57.0 47.0 58.8 51.2 52.0 58.7 32.8 39.3 45.7
ViLAMP 41.4 29.3 48.8 31.2 37.7 59.8 62.4 45.3 20.0 46.9 55.7 49.7 52.0 36.7 48.5 53.3 51.0 32.3 31.6 42.1
VideoLLaMA 30.8 21.6 11.5 20.5 21.1 28.4 23.0 23.9 20.0 23.8 22.9 41.1 23.2 15.4 25.7 40.1 26.6 33.1 26.2 31.5
VideoLLaMA2.1 38.0 30.6 47.0 30.4 36.5 34.6 52.7 41.5 16.7 36.4 53.8 60.3 27.0 33.3 43.6 44.0 31.6 32.1 23.7 32.9
MIO 29.3 20.3 29.3 10.7 22.4 4.5 18.2 10.4 23.3 14.1 22.9 39.7 29.0 19.2 27.7 42.1 31.6 16.1 19.6 27.4
ImageBind-LLM 21.3 22.4 25.0 11.6 20.1 23.5 13.3 23.8 19.5 20.0 15.4 32.4 24.0 23.3 23.8 45.0 25.1 28.9 22.9 30.5
ChatBridge 28.9 18.0 30.2 16.1 23.3 31.4 30.9 23.7 8.7 23.7 24.1 43.7 41.0 16.9 31.4 43.7 25.3 30.4 24.4 31.0
OneLLM 27.0 26.1 34.1 7.1 23.6 29.0 36.4 20.6 27.8 28.5 24.9 49.0 23.0 22.4 29.8 43.4 26.4 29.5 23.2 30.6
Qwen2.5-Omni 42.2 30.1 36.6 33.0 35.5 56.4 64.2 41.5 15.8 44.5 45.8 55.0 28.0 24.3 38.3 71.1 48.4 27.0 71.8 54.6
GPT-4o (20241120) 36.5 43.6 25.6 28.5 33.6 49.2 77.0 47.2 32.5 50.9 62.1 62.9 52.0 71.2 62.1 - - - - -
GPT-5 (20250807) 43.7 48.9 59.8 34.8 46.8 69.8 84.8 64.2 58.3 69.5 73.1 66.9 64.0 65.0 67.3 - - - - -
Gemini-1.5-Pro 49.4 51.9 53.0 49.1 50.9 87.1 73.9 52.8 71.7 71.4 53.4 60.3 54.0 75.1 60.7 90.1 76.8 66.4 84.6 79.5
Gemini-2.5-flash 52.1 57.1 59.1 43.8 53.0 92.2 77.6 61.3 71.7 75.7 65.6 68.9 58.0 67.2 64.9 98.7 80.6 57.7 99.1 84.0
Gemini-2.5-Pro 54.4 54.1 53.0 50.0 52.9 95.0 84.8 69.8 84.2 83.5 79.0 72.2 63.0 68.4 70.7 96.7 78.7 72.3 98.3 86.5
Human 87.6 85.0 87.8 78.0 84.6 98.9 84.1 92.5 78.3 88.5 86.7 84.5 88.6 88.1 87.0 96.0 96.1 87.0 98.3 94.4
Figure 7: Histogram of smoothed distractor selection frequency by 25 open-source MLLMs across six tasks, generated via Distractor-Included Synthesis QA pipeline.
Appendix E Distractor Quality Analysis
To quantitatively assess the quality of the distractors generated by our synthesis pipeline, we compute the selection frequency of every distractor across the 25 open-source MLLMs evaluated in our study, and visualize the resulting histogram distribution (Figure E). The analysis reveals two distinctive properties. First, the distractors exhibit exceptionally high effectiveness: the number of items at X=0 is extremely small, and on six tasks, approximately 90% of distractors are selected by at least one model. This demonstrates that our “model-induced error” generation strategy successfully avoids trivial or obviously incorrect options, compelling models to perform fine-grained reasoning over all choices. Second, the distribution forms an ideal discriminatory pattern, with the majority of distractors selected by 1 to 8 models and a clear long-tail structure. This indicates that most distractors effectively attract and differentiate about 4%–32% of the models, reflecting high-quality deceptiveness and strong discriminatory power. The long-tail pattern further suggests a healthy difficulty gradient: the benchmark neither becomes too easy (which would render distractors ineffective) nor too difficult (which would cause random guessing).
Overall, the combination of a low-frequency peak and a long-tailed distribution provides strong evidence that our distractors are both well-formed and discriminative, forming a solid foundation for a reliable evaluation benchmark.
Moreover, model accuracy results strongly validate this effect: performance follows the expected capability hierarchy—commercial frontier models (e.g., Gemini-2.5-Pro, Gemini-1.5-Pro) outperform advanced open-source 7B models (e.g., Qwen3-VL, InternVL3.5), which in turn surpass earlier models of similar scale (e.g., Video-LLaVA, ShareGPT4Video). This gradient confirms that the distractors successfully penalize weaker models while stronger models remain more resilient, further showcasing the benchmark’s discriminatory strength and reliability.
Appendix F Modality Ablation in VideoLLaMA2
@@
Models Input Modal Speech-Visual Alignment
AVSM ASD AVAD SCM Avg
Random 42.8 23.6 33.3 25.0 31.2
Video-LLaMA2.1-AV A, V 44.0 31.6 32.1 23.7 32.9
Video-LLaMA2.1-AV V 43.4 29.0 32.8 18.8 31.0
VideoLLaMA2-7B-16F V 47.4 38.1 32.1 15.4 33.3
@@
Appendix G Impact of Frame Sampling
Table 6: Impact of different frame sampling strategies on HUMANVBENCH. We report the accuracy (%) and total inference time (minutes) for InternVL3.5 and LLaVA-Video. Denser sampling beyond 8 frames leads to diminishing returns and increased computational overhead.
Model InternVL3.5 LLaVA-Video
Sampling Rate 4f 8f 16f 4f 8f 16f
Emotion (4 tasks) 45.2 46.1 47.2 36.9 39.4 38.8
Person Recog. (4 tasks) 55.0 57.6 56.5 52.2 53.2 53.9
Behavior (4 tasks) 57.0 58.6 59.9 51.3 52.0 53.0
AV-Sync (4 tasks) 44.8 46.0 45.8 42.3 44.0 43.5
Overall Avg. 50.5 52.1 52.3 45.7 47.2 47.3
Total Time (min.) 28.3 39.2 62.4 28.7 40.9 62.3
To investigate the effect of temporal resolution on HUMANVBENCH, we evaluate representative MLLMs (InternVL3.5 and LLaVA-Video) across 4, 8, and 16 frames. As shown in Table 6, while denser sampling generally yields performance gains, we observe significant diminishing returns. Specifically, for InternVL3.5, increasing from 8 to 16 frames results in only a marginal $0.26\%$ overall improvement while incurring a $1.6\times$ increase in latency (from 39.2 to 62.4 minutes). Notably, in tasks like Person Recognition, 8-frame sampling even outperforms 16-frame sampling, suggesting that excessive frames may introduce redundant temporal noise. These findings underscore the importance of efficient temporal modeling over brute-force dense sampling.
Appendix H Impact of Human Density on Model Performance
Table 7: Performance across varying human densities. Human density is categorized as Low (1–3 persons), Med (4–8), and High (>8>8). Accuracy (%) is reported for Strong and Mid-tier groups.
Task Category Strong Models Mid-tier Models
(Human Density) Low Med High Low Med High
Overall Average 63.5 57.2 47.2 37.3 36.8 35.6
Emotion Perception 49.0 50.8 50.0 35.5 34.9 39.5
Human Recognition 78.1 65.1 36.1 42.4 38.0 31.9
Behavior Analysis 61.6 56.0 55.6 33.8 40.2 35.3
Speech Alignment 65.1 57.0 37.3 34.2
We evaluate two model tiers—Strong (InternVL3.5, Qwen2-VL-72B, VideoLLaMA3) and Mid-tier (InternVL2, CogVLM2-Video, Qwen2-VL-7B)—to benchmark robustness against varying human densities. As shown in Table 7, a capability collapse occurs in autonomous localization (e.g., Human Recognition), where Strong models’ accuracy plunges from 78.1% to 36.1% in dense scenes. Conversely, performance in visually-grounded tasks (e.g., Emotion Perception, where targets are pre-marked by bounding boxes) remains stable. This indicates that multi-entity interference primarily hinders autonomous spatial-temporal grounding—the ability to isolate a specific individual from a crowd—rather than the high-level reasoning itself.
While Strong models maintain a performance lead in all scenarios, this margin narrows in dense environments; for instance, the Overall Average gap between the two tiers shrinks from 26.2% to 11.6%. This underscores the need for density-invariant grounding to overcome interference in crowded scenes. Future work should prioritize higher-resolution attention mechanisms to maintain precise focus on specific targets amidst increasing environmental complexity.
Appendix I Model Evaluation Implementation
Prompt. In order to facilitate the statistical model to answer the results, following common practices used in MLLM evaluations, we adopt the following prompt to guide the MLLM to output option letters: “ Select the best answer to the following multiple-choice question based on the video. Respond with only the letter of the correct option. <Question-choices>   Only answer best answer’s option letter. Your option is: ”. Evaluation Environments. All evaluation experiments for open-source models were conducted on a single NVIDIA L20 GPU with an inference batch size of 1.
@@ Appendix I Model Evaluation Implementation
Model Time (min) top_p top_k num_beams temp. frames
Chat-UniVi (7B) 40 1 50 1 0.2 1 f/s
CogVLM2-Video (8B) 37 0.1 1 1 0.2 1 f/s
Video-LLaVA (7B) 49 1 50 1 1 8 f
LLaVA-OneVision (7B) 49 1 50 1 1 8 f
PLLaVA (7B) 41 0.9 50 1 1 16 f
ShareGPT4Video (8B) 50 0.9 50 1 1 16 f
Otter-V (7B) 62 1 50 3 1 16 f
VideoChat2-IT (7B) 53 0.9 50 1 1 16 f
InternVL2 (7B) 44 1 52 1 1.0 8 f
InternVL2.5 (7B) 31 1 52 1 1.0 8 f
InternVL3 (7B) 186 1 52 1 1.0 64 f
InternVL3.5 (7B) 107 1 52 1 1.0 32 f
Qwen2-VL (7B) 45 0.001 1 1 0.1 2 f/s
Qwen2.5-VL (7B) 35 0.001 1 1 0.1 2 f/s
Qwen3-VL (7B) 34 0.001 1 1 0.1 2 f/s
LLaVA-Video (7B) 182 0.8 20 1 0.7 64 f
Video-LLaMA3 (7B) 90 0.8 20 1 0.7 1 f/s
InternVideo2.5 (7B) 170 1 50 1 1 128f
ViLAMP (7B) 29 0.8 20 1 0.7 1 f/s
Video-LLaMA (7B) 40 1 50 2 1 8 f
Video-LLaMA2.1 (7B) 31 0.9 50 1 0.2 8 f
ImageBind-LLM (7B) 77 1 50 1 1 15 f
ChatBridge (13B) 29 1 50 1 0.2 4 f
OneLLM (7B) 130 0.75 50 1 0.1 15 f
MIO (7B) 88 0.7 0 1 1 FPS/10
Qwen2.5_Omni (7B) 91 1 50 1 1 2 f/s
GPT-4o (20241120) 191 1 - - 1 FPS/10
Gemini-1.5-Pro 254 1 40 - 0.9 API
@@
Appendix J Definitions and Examples for Each Task
J.1 Emotion Perception
@@
Figure 11: Example of Emotion Intensity Comparision task.
J.2 Person Recognition
@@
J.3 Human Behavior Analysis
@@
J.4 Cross-Modal Speech-Visual Alignment
@@
Appendix K Annotations Details and Examples in Human-Centric Annotation Pipeline
@@
In the video_facial_description_mapper, we use the face bounding box track to crop the video, creating face-focused reconstructed videos for emotion description using Qwen2.5-VL [15].
@@
Appendix L Complete Construction Details of All Tasks
@@
L.1 Construction Details of 6 Descriptive Human-Centric Questions
We first present the general instruction templates in the six task generation processes.
The prompt template for the Video-MLLMs used to obtain an answer based on the question is: Answer the following question based on the video. Respond in one concise sentence. ⟨question_text⟩. The obtained answer may later serve as the ensemble best answer or as material for distractors.
The prompt used to rank all candidate answers is:
Task: Given the following question and its four answer options, rank the options in order of accuracy from highest to lowest. Question: ⟨question⟩, Options: answer1: ⟨answers[0]⟩, answer2: ⟨answers[1]⟩, answer3: ⟨answers[2]⟩, answer4: ⟨answers[3]⟩. Please return the ranking in the following format: [(“answer1”, “Answer content”, Accuracy score), (“answer2”, “Answer content”, Accuracy score), (“answer3”, “Answer content”, Accuracy score), (“answer4”, “Answer content”, Accuracy score)]. We then aggregate the accuracy scores of each answer across all MLLMs. Accounting for model capability differences, we assign different weights; in our setup, Gemini’s rankings receive twice the weight of the others. The answer with the highest total score is treated as the temporary correct option, while the remaining answers serve as material for constructing distractors.
@@
L.2 Construction Details of 10 Closed-Ended Human-Centric Questions
@@
L.3 Details of Human Efforts for HumanVBench
@@
Manual Verification and Correction: Except for the tasks that are already reliable enough, which include three tasks derived from the aforementioned human annotations, the Audio-Visual Alignment Detection and Speech Content Matching tasks, all other tasks require manual verification to ensure quality, following the process described in Section 3.3. During correction, low-quality samples (e.g., person transitions in human tracking, video freezing midway) are required to be flagged for removal.